Obtain the least-order differential equation x = c1sin (y + c2)?


1
Expert's answer
2020-12-04T13:47:16-0500

"x=c_1\sin(y+c_2)\\
\text{differentiate both sides with respect to y}\\
x'=c_1\cos(y+c_2)\\
\text{differentiate both sides with respect to y again}\\
x''=-c_1\sin(y+c_2)\\
x''=-x\\
x''+x=0."